2. The coordinates of the vertices of ΔARC are A(3,3), R(1,-1), and C(-2,1). Angela is trying to determine whether this shape is a right triangle or not. She has made a mistake in the work below the graph. Explain the mistake and show the correct solution.

Hint: The grid is provided for your use (optional). The graph is not worth any points and will not be considered when graded.

Hint: When explaining the mistake, make sure you identify how it is wrong AND how it needs to be corrected.


(12 points total)

 Angela’s solution: 

slope of AR = 

slope of RC = 

slope of AC = 

Since AR and RC are opposite reciprocal slopes, those two sides are perpendicular, and the triangle is a right triangle.




Answer:






1
Expert's answer
2020-10-25T18:54:27-0400

"\textsf{Denote the slope of the line segment}\\
\textsf{joining points}\, A \, \textsf{and}\, R \, \textsf{by}\, AR_s,\\
A \, \textsf{and}\, C \, \textsf{by}\, AC_s \, \textsf{and}\, R \, \textsf{and}\, C \, \textsf{by}\, RC_s\\

\begin{aligned}
AR_s &= \frac{3 - (-1)}{3 - 1} = \frac{4}{2} = 2\\
AC_s &= \frac{3 - 1}{3 - (-2)} = \frac{2}{5}\\
RC_s &= \frac{1 - (-1)}{-2 - 1} = \frac{-2}{3}
\end{aligned}\\

\textsf{By the condition of perpendiicularity, if the slope}\\
\textsf{of two lines are}\, m_1 \, \textsf{and}\, m_2,\\
\textsf{they are said to be perpendicular if}\, m_1 m_2 = -1\\

\textsf{The mistake was that Angela mistakened} -3 \, \textsf{for}\, -2 \\
\textsf{in calculating the slope of line}\, RC \\

\textsf{If}\, -2 \, \textsf{was replaced with}\, -3, RC_s = \frac{-2}{4} = \frac{-1}{2}\\

RC_s \, \textsf{would have been}\, \frac{-1}{2} = \frac{-1}{AR_s} \\

\textsf{and as such,}\, RC \, \textsf{would be perpendicular to}\,AR \\
\textsf{making the triangle a right-angled triangle.}\\

\textsf{For triangle} \, ARC \, \textsf{to be a right-angled triangle},\\
\textsf{one of the line segments joining the points}\\
\textsf{has to be perpendicular to the other.}\\


AR_s RC_s \neq -1, AR_s AC_s \neq -1,\, \&\, AC_s RC_s \neq -1 \\

\textsf{Therefore, since none of the slopes are opposite}\\
\textsf{reciprocal of the other, the triangle is}\\
\textsf{not a right-angled triangle.}"